# Current address: ]College of William and Mary, Williamsburg, Virginia 23187-8795

### v2

@@ I Introduction
The principle of $\chi$EFT is that it describes the nucleon in terms of effective hadronic degrees of freedom instead of the fundamental quark and gluon fields. It is completely non-perturbative in terms of $\alpha_{s}$ because it employs series whose expansion parameters are based e.g. on the pion mass, the nucleon-$\Delta_{1232}$ mass gap or the characteristic energy scale of the chiral symmetry breaking. However, while leading-order (LO) calculations of $\chi$EFT are relatively tractable, next-to-leading-order (NLO) and higher order calculations are much more involved. An important experimental task is therefore to provide precise hadron structure data that can test, and potentially guide, $\chi$EFT calculations and other approaches to non-perturbative QCD. Such data must cover the domain of low momentum transfer squared ($Q^{2}$) where $\chi$EFT applies, typically below a few tenths of GeV2. Especially important in this context are low $Q^{2}$ nucleon spin structure data [10, 11, 11, 12, 13, 14, 15, 16, 17, 18] which are challenging to obtain experimentally but are crucial for the development of $\chi$EFT [19]. One such experimental endeavor is that of the Jefferson Lab (JLab) Experimental Group EG4, constituted of experiments E03006 and E05111 that used the CEBAF Large Acceptance Spectrometer (CLAS) in Hall B [20]. EG4 was designed to test $\chi$EFT predictions by precisely measuring the longitudinal spin structure function $g_{1}$ and its moments for the proton and the deuteuron down to $Q^{2}\simeq 0.01$ GeV2. These low $Q^{2}$ data, reported in two letters [21, 22], provided benchmarks for $\chi$EFT calculations. This document is the archival article that presents in detail the EG4 experiments, data analysis procedure for the inclusive scattering channel, results on the proton and the deuteron, and followup analysis and results on the neutron spin structure functions. Additional results on the target- and double-spin asymmetries of pion electroproduction are available [23], but these were focused on the semi-inclusive channel and are not included here.
@@
The article is arranged as follows: In Section II we present the formalism of unpolarized and polarized electron scattering. We introduce the nucleon structure functions, their moments, and polarized sum rules, which are relevant not only for testing $\chi$EFT but also to form observables that have no known method of direct measurement. In Section III we describe EG4 and the special features that distinguish it from other CLAS experiments. In Section IV we describe the analysis procedure of extracting the polarized yield differences from the data. In Section V we describe the Monte-Carlo simulation, the normalization of the data using the elastic scattering yield, and the extraction of the structure function $g_{1}$. In Section VI.1 we present results on $g_{1}$, $A_{1}F_{1}$ of the proton and the deuteron, followed by the procedure to extract the neutron $g_{1}$ results. We present results for and discussions of all the moments in Section-VII. We then summarize, conclude, and provide perspectives in Section VIII.
@@ II.4.2 Gerasimov-Drell-Hearn (GDH) Sum Rule
$\nu_{thr}$ $=$ $m_{\pi}+\frac{m_{\pi}^{2}+Q^{2}}{2M_{T}},~{}$ (28)
with $m_{\pi}$ the pion mass. The anomalous magnetic moment of the target $\kappa$ is defined by the total magnetic moment of the particle $\bm{\mu}=\frac{e(Q/e+\kappa)}{M_{T}}\bm{S}$ with $Q/e$ the charge of the target in units of the elementary charge, and can be related to the gyromagnetic ratio $g$ as $\kappa=g/2-1$. The RHS of Eq. (27) is $-204~{}\mu$b for the proton ($S=\frac{1}{2}$, $\kappa=1.793$ and, $g/2=2.793$), $-234~{}\mu$b for the neutron ($S=\frac{1}{2}$, $\kappa=-1.913$, and $g/2=-1.913$), and $0.65~{}\mu$b for the deuteron ($S=1$, $\kappa=-0.143$, and $g/2=0.857$). Whereas the overall consensus [63] is that the GDH sum rule is theoretically very solid, the question of its validity has been debated, mainly regarding whether $\sigma_{TT}/\nu$ decreases fast enough with $\nu$ for the integral to converge. Concurrently, its experimental verification has been the focus of several experiments at MAMI and ELSA [65, 66, 67, 68, 69], BNL [70], JLab [21, 21, 71], GRAAL (ESRF) [72], and HIGS (TUNL) [73] providing results that show, given reasonable assumptions for the large $\nu$ behavior of its integrand, the GDH sum rule is valid for the proton to within the 10% precision of the world data [63, 21, 22, 74].
@@ II.5 Deuteron Nuclear Model
For moments of spin structure functions, the most conspicuous nuclear effect, viz. smearing of the resonances by Fermi motion, is integrated out. It vanishes exactly for first moments and to a good approximation for higher moments. We studied this assertion using a model for both smeared and unsmeared spin structure functions and found no difference in the first moments, and only minute differences in the higher moments ($\overline{\gamma}_{0}$) (up to 4.8%, which is small compared to our statistical and systematic uncertainties). No corrections were applied to the data; we only account for the effective polarization of the nucleons in deuterium when relating integrals for protons, neutrons, and deuterons, while supplementing the region 1.073 GeV $<W<$ 1.15 GeV with the MAID parametrization:
@@ III.2 The Polarized Targets
During EG4 the polarized target was placed 1.01 m upstream from the CLAS center to increase the acceptance at low $Q^{2}$ by reducing the minimum angle for scattered electrons. The targets included frozen 15NH3 and 15ND3 dynamically polarized at 1 K with a 5-T field, as well as two carbon targets and an empty cell for reference. These were the same as the targets used for previous CLAS double-polarization measurements [125]. The beam was rastered to protect the target from overheating and to minimize depolarization. The target density was 1.056 and 0.917 g/cm3 for NH3 and ND3, respectively.
@@ IV.3 Particle Identification and Cherenkov Efficiency
$\eta=\frac{N_{2}}{N_{1}}~{},$ $\Delta\eta_{N_{2}\neq N_{1}}\sqrt{\frac{N_{2}(1-N_{2}/N_{1})}{N_{1}}}~{}.$ (40)
@@ V.3 Smearing Factor for DC Resolution
By comparing the $W$ peaks of the elastic simulation those obtained from data, it was found necessary to apply DC smearing in the GPP stage of the simulation. For the ND3 analysis, a single DC smearing factor for each beam energy was used. For the NH3 analysis, we found it necessary to adjust the smearing factor separately for each beam energy. This was done by adopting a wire-number dependence in the DC smearing factor of GPP. The functional form used for the DC smearing factor $y$ is
@@ V.6 Extraction of $g_{1}$ and $A_{1}F_{1}$
  • The choice for $F_{1,2}$ and $R$ was based on the latest fit by Bosted/Christie/Kalantarians (dated 2014), interpolating to the real photon point at very low $Q^{2}$. The most up-to-date references are [141] and [142].
@@ VII.3 Moments of the Proton Spin Structure functions
Figure 18: [Color online] EG4 results on the proton ${\overline{\gamma}_{0}}^{p}(Q^{2})$. Across all $Q^{2}$ bins, the size of the systematic uncertainty (red colored band) varies from $0.0045$ to $0.198$, while the relative size of the systematic to statistical uncertainty varies within $(66-364)\%$. The photoproduction data point [153] is slightly displaced from $Q^{2}=0$ (purple star). Note that phenomenological model predictions [54, 56] are not available for this integral. See Fig. 16 caption for more details.
@@ VII.4 Moments of the Deuteron Spin Structure Functions
Figure 21: [Color online] EG4 results on the deuteron $\overline{\gamma}_{0}^{d}(Q^{2})$. Across all $Q^{2}$ bins, the size of the systematic uncertainty (red colored band) varies from $0.0123$ to $0.339$, while the relative size of the systematic to statistical uncertainty varies within $(37-93)\%$. Note that phenomenological model predictions [54, 56] are not available for this integral. See Fig. 16 caption for more details.
@@ VII.5 Moments of the Neutron Spin Structure Functions
Figure 22: [Color online] EG4 results on the neutron $\overline{\Gamma}_{1}^{n}(Q^{2})$, compared with earlier results from EG1b (d,p) [102, 103] (black crossbars), JLab E97110 (neutron extracted from 3He) [71] (blue triangles), and E94010 (3He) [93] (green triangles). The size of the systematic uncertainty (red colored band) varies from $0.0013$ to $0.0071$, while the relative size of the systematic to statistical uncertainty varies from $(45-92)\%$, from the lowest to the highest $Q^{2}$ bins. See Fig. 16 caption for more details.
Figure 23: [Color online] EG4 results on the neutron $\overline{I}_{TT}^{n}(Q^{2})$, compared with results from JLab E94010 (3He) [92] (green triangles) and E97110 (3He) [71] (blue triangles). Across all $Q^{2}$ bins, the size of the systematic uncertainty varies from $0.0275$ to $0.141$, while the relative size of the systematic to statistical uncertainty varies within $(49-123)\%$. The GDH value is shown as the short horizontal line at $\overline{I}^{n}_{TT}(0)=-0.915$. Note that earlier CLAS data [102] and the phenomenological model predictions [54, 56] are not available for this integral. See Fig. 16 caption for more details.
The resulting $\overline{\Gamma}_{1}^{n}(Q^{2})$ and $\overline{I}_{TT}^{n}(Q^{2})$ are shown in Figs. 22 and 23, respectively. They agree with previous data from E143 [107], EG1a [94, 95] and EG1b [102, 103] where the neutron information was extracted from the combined deuteron and proton data, as well as those from E94010 [93] and E97110 [71] for which the neutron information was extracted from 3He data. There is also agreement with the phenomenological models of Burkert and Ioffe [53, 54] and Pasechnik et al. [56] (when available) and with the $\chi$EFT predictions from Bernard et al. [59] (up to $Q^{2}\simeq 0.08$ GeV2) and, to a lesser extent, with Alarcón et al. [60]. The MAID model prediction is systematically more positive than the data. Like in the case of $\overline{\gamma}_{0}^{n}$, the calculation of [154] (not shown in the figure) agrees in sign and behavior with the $\overline{\gamma}_{0}^{n}(Q^{2})$, up to a similar normalization factor as for $\overline{\gamma}_{0}^{p}(Q^{2})$.
@@
Results on the generalized longitudinal spin polarizability on the neutron are shown in Fig. 24, along with earlier results from JLab EG1b [102, 103]), E97110 [71] and E94010 [93]). The EG4 neutron results generally agree with the earlier data, though there is a tension between EG4 and E97110 for $Q^{2}\lesssim 0.1$ GeV2. Adding the EG4 deuteron and proton systematic uncertainties linearly, rather than in quadrature, does not remove the tension. It contrasts with the agreement between EG4 and E97110 for the neutron data on $\overline{\Gamma}_{1}^{n}$ or $\overline{I}_{TT}^{n}$, suggesting the discrepancy arises from the high-$x$ (near pion threshold) contribution. The EG4 results agree with $\chi$EFT calculation from Ref. [59] up to $Q^{2}\approx 0.1$ GeV2 but not with that of Alarcón et al. [60].
Figure 24: [Color online] EG4 results on the generalized spin polarizability $\overline{\gamma}_{0}^{n}$, compared with earlier results from JLab EG1b (d,p) [102, 103] (black crossbars), E97110 (3He) [71] (blue triangles), and E94010 (3He) [93] (green triangles). Across all $Q^{2}$ bins, the size of the systematic uncertainty varies from $0.0158$ to $0.397$, while the relative size of the systematic to statistical uncertainty varies from $(40-124)\%$. Note that phenomenological model predictions [54, 56] are not available for this integral. See Fig. 16 caption for more details.
@@ VII.6 Isospin Analysis on the Moments of Spin Structure Functions
where $\omega_{D}$ is given under Eq. (38) and with the deuteron moments understood as “per nucleus”. Alternatively to using the EG4 proton and deuteron data, one can use the proton results from EG4 and the neutron results obtained from 3He data such as those from JLab E97110 [71] to perform the isospin separation.
@@ VII.6.1 Results on the Bjorken Sum $\overline{\Gamma}_{1}^{p-n}(Q^{2})$
Figure 25: [Color online] The Bjorken sum $\overline{\Gamma}_{1}^{p-n}(Q^{2})$ from EG4 (solid red circles) and from the EG4 proton data combined with the JLab E97110 neutron data [71] (blue triangles). Also shown are earlier data from E94010/EG1a [42] (green triangles), CLAS EG1b [43] (black crossbars), and EG1dvcs [44] (brown open squares). The inner (outer) error bars on the data points give their statistical (total) uncertainties, while the systematic uncertainties of EG4 and EG4/E97110 are also given by the red and the blue bands, respectively. Across all $Q^{2}$ bins, the size of the systematic uncertainty of the EG4 results (red colored band) varies from $0.0017$ to $0.0115$, while the relative size of the systematic to statistical uncertainty varies from $(58-131)\%$. The $\chi$EFT predictions from Alarcón et al. [60] and Bernard et al. [59] (the latter without error band as given in Ref. [59]) are shown along with phenomenological model predictions from Burkert-Ioffe [53, 54], Pasechnik et al. [56], MAID [113], HLFQCD [58], and Gabdrakhmanov et al. [161].
@@
Note the comparison between the $\overline{\Gamma}_{1}^{p-n}$ results obtained from EG4’s $p$ and D data and those extracted from the EG4 $p$ and E97-110 $n$(3He) data. An agreement between the two different approaches was not assured since the nuclear structures of the deuteron and 3He are quite different: 3He is strongly bound compared to the deuteron, therefore more sensitive to potential nuclear modification of the proton. Our treatment of the deuteron is minimal and consists of accounting for its $D$-state [116, 117, 118, 119, 120], while treatment of the 3He data accounted for the nucleon effective polarizations [162]. As shown in Fig. 25, the two approaches agree well with each other. This suggests that, at least for $\overline{\Gamma}_{1}^{p-n}$, the nuclear corrections applied to the D and 3He data to obtain the neutron information appear sufficient at these low $Q^{2}$ values.
@@
Both the EG4 and EG4/E97110 results on $\overline{\Gamma}_{1}^{p-n}$ are systematically above $\chi$EFT and model predictions, except that of HLFQCD [58] and [161], which agree with the data. To allow a quantitative comparison between the data and theory, we employed a fit of the form $bQ^{2}+cQ^{4}$ to the entire world data set, based on Eq. (32[52], where $b$ is expected to be the GDH slope. Our results are shown in Table 4. The “$uncor$” uncertainty designates the point-to-point uncorrelated uncertainty. It is the quadratic sum of the statistical uncertainty and a fraction of the systematic uncertainty determined so that $\chi^{2}$/n.d.f = 1 for the best fit. The “$cor$” uncertainty is the correlated uncertainty estimated from the remaining fraction of the systematic uncertainty.
[…]

### v3

# Current address: ]College of William and Mary, Williamsburg, Virginia 23187, USA

@@ I Introduction
The principle of $\chi$EFT is that it describes the nucleon in terms of effective hadronic degrees of freedom instead of the fundamental quark and gluon fields. It is completely non-perturbative in terms of $\alpha_{s}$ because it employs series whose expansion parameters are based e.g. on the pion mass, the nucleon-$\Delta_{1232}$ mass gap or the characteristic energy scale of the chiral symmetry breaking. However, while leading-order (LO) calculations of $\chi$EFT are relatively tractable, next-to-leading-order (NLO) and higher order calculations are much more involved. An important experimental task is therefore to provide precise hadron structure data that can test, and potentially guide, $\chi$EFT calculations and other approaches to non-perturbative QCD. Such data must cover the domain of low momentum transfer squared ($Q^{2}$) where $\chi$EFT applies, typically below a few tenths of GeV2. Especially important in this context are low $Q^{2}$ nucleon spin structure data [10, 11, 11, 12, 13, 14, 15, 16, 17, 18] which are challenging to obtain experimentally but are crucial for the development of $\chi$EFT [19]. One such experimental endeavor is that of the Jefferson Lab (JLab) Experimental Group EG4, constituted of experiments E03-006 and E05-111 that used the CEBAF Large Acceptance Spectrometer (CLAS) in Hall B [20]. EG4 was designed to test $\chi$EFT predictions by precisely measuring the longitudinal spin structure function $g_{1}$ and its moments for the proton and the deuteuron down to $Q^{2}\simeq 0.01$ GeV2. These low $Q^{2}$ data, reported in two letters [21, 22], provided benchmarks for $\chi$EFT calculations. This document is the archival article that presents in detail the EG4 experiments, data analysis procedure for the inclusive scattering channel, results on the proton and the deuteron, and followup analysis and results on the neutron spin structure functions. Additional results on the target- and double-spin asymmetries of pion electroproduction are available [23], but these were focused on the semi-inclusive channel and are not included here.
@@
The article is arranged as follows: In Section II we present the formalism of unpolarized and polarized electron scattering. We introduce the nucleon structure functions, their moments, and polarized sum rules, which are relevant not only for testing $\chi$EFT but also to form observables that have no known method of direct measurement. In Section III we describe EG4 and the special features that distinguish it from other CLAS experiments. In Section IV we describe the analysis procedure of extracting the polarized yield differences from the data. In Section V we describe the Monte-Carlo simulation, the normalization of the data using the elastic scattering yield, and the extraction of the structure function $g_{1}$. In Section VI.1 we present results on $g_{1}$, $A_{1}F_{1}$ of the proton and the deuteron, followed by the procedure to extract the neutron $g_{1}$ results in Section VI.2. We present results for and discussions of all the moments in Section-VII. We then summarize, conclude, and provide perspectives in Section VIII.
@@ II.4.2 Gerasimov-Drell-Hearn (GDH) Sum Rule
$\nu_{thr}$ $=$ $m_{\pi}+\frac{m_{\pi}^{2}+Q^{2}}{2M},~{}$ (28)
with $m_{\pi}$ the pion mass. The anomalous magnetic moment of the target $\kappa$ is defined by the total magnetic moment of the particle $\bm{\mu}=\frac{e(Q/e+\kappa)}{M_{T}}\bm{S}$ with $Q/e$ the charge of the target in units of the elementary charge, and can be related to the gyromagnetic ratio $g$ as $\kappa=g/2-Q/e$. The RHS of Eq. (27) is $-204~{}\mu$b for the proton ($S=\frac{1}{2}$, $\kappa=1.793$ and, $g/2=2.793$), $-234~{}\mu$b for the neutron ($S=\frac{1}{2}$, $\kappa=-1.913$, and $g/2=-1.913$), and $0.65~{}\mu$b for the deuteron ($S=1$, $\kappa=-0.143$, and $g/2=0.857$). Whereas the overall consensus [63] is that the GDH sum rule is theoretically very solid, the question of its validity has been debated, mainly regarding whether $\sigma_{TT}/\nu$ decreases fast enough with $\nu$ for the integral to converge. Concurrently, its experimental verification has been the focus of several experiments at MAMI and ELSA [65, 66, 67, 68, 69], BNL [70], JLab [21, 21, 71], GRAAL (ESRF) [72], and HIGS (TUNL) [73] providing results that show, given reasonable assumptions for the large $\nu$ behavior of its integrand, the GDH sum rule is valid for the proton to within the 10% precision of the world data [63, 21, 22, 74].
@@ II.5 Deuteron Nuclear Model
For moments of spin structure functions, the most conspicuous nuclear effect, viz. smearing of the resonances by Fermi motion, is integrated out. It vanishes exactly for first moments and to a good approximation for higher moments. We studied this assertion using a model for both smeared and unsmeared spin structure functions and found no difference in the first moments, and only minute differences in the higher moments ($\overline{\gamma}_{0}$) (up to 4.8%, which is small compared to our statistical and systematic uncertainties). Thus, no corrections were applied to the data; we only account for the effective polarization of the nucleons in deuterium when relating integrals for protons, neutrons, and deuterons, while supplementing the region 1.073 GeV $<W<$ 1.15 GeV with the MAID parametrization:
@@ III.2 The Polarized Targets
During EG4 the polarized target was placed 1.01 m upstream from the CLAS center to increase the acceptance at low $Q^{2}$ by reducing the minimum angle for scattered electrons. The targets included frozen 15NH3 and 15ND3 dynamically polarized at 1 K with a 5-T field, as well as two carbon targets and an empty cell for reference. These were the same as the targets used for previous CLAS double-polarization measurements [125]. The beam was rastered to protect the target from overheating and to minimize depolarization. The target density was 0.917 and 1.056 g/cm3 for NH3 and ND3, respectively.
@@ IV.3 Particle Identification and Cherenkov Efficiency
$\eta=\frac{N_{2}}{N_{1}}~{},$ $\Delta\eta_{N_{2}\neq N_{1}}=\sqrt{\frac{N_{2}(1-N_{2}/N_{1})}{N_{1}}}~{}.$ (40)
@@ V.3 Smearing Factor for DC Resolution
By comparing the $W$ peaks of the elastic simulation with those obtained from data, it was found necessary to apply DC smearing in the GPP stage of the simulation. For the ND3 analysis, a single DC smearing factor for each beam energy was used. For the NH3 analysis, we found it necessary to adjust the smearing factor separately for each beam energy. This was done by adopting a wire-number dependence in the DC smearing factor of GPP. The functional form used for the DC smearing factor $y$ is
@@ V.6 Extraction of $g_{1}$ and $A_{1}F_{1}$
  • The choice for $F_{1,2}$ and $R$ was based on the latest fit by Bosted/Christy/Kalantarians (dated 2014), interpolating to the real photon point at very low $Q^{2}$. The most up-to-date references are [141] and [142].
@@ VII.3 Moments of the Proton Spin Structure functions
Figure 18: [Color online] EG4 results on the proton ${\overline{\gamma}_{0}}^{p}(Q^{2})$. Across all $Q^{2}$ bins, the size of the systematic uncertainty (red colored band) varies from $0.0045$ to $0.198$ (×$\times 10^{-4}$ fm4), while the relative size of the systematic to statistical uncertainty varies within $(66-364)\%$. The photoproduction data point [153] is slightly displaced from $Q^{2}=0$ (purple star). Note that phenomenological model predictions [54, 56] are not available for this integral. See Fig. 16 caption for more details.
@@ VII.4 Moments of the Deuteron Spin Structure Functions
Figure 21: [Color online] EG4 results on the deuteron $\overline{\gamma}_{0}^{d}(Q^{2})$. Across all $Q^{2}$ bins, the size of the systematic uncertainty (red colored band) varies from $0.0123$ to $0.339$ (×$\times 10^{-4}$ fm4), while the relative size of the systematic to statistical uncertainty varies within $(37-93)\%$. Note that phenomenological model predictions [54, 56] are not available for this integral. See Fig. 16 caption for more details.
@@ VII.5 Moments of the Neutron Spin Structure Functions
Figure 22: [Color online] EG4 results on the neutron $\overline{\Gamma}_{1}^{n}(Q^{2})$, compared with earlier results from EG1b (d,p) [102, 103] (black crossbars), JLab E97-110 (neutron extracted from 3He) [71] (blue triangles), and E94-010 (3He) [93] (green triangles). The size of the systematic uncertainty (red colored band) varies from $0.0013$ to $0.0071$, while the relative size of the systematic to statistical uncertainty varies from $(45-92)\%$, from the lowest to the highest $Q^{2}$ bins. See Fig. 16 caption for more details.
Figure 23: [Color online] EG4 results on the neutron $\overline{I}_{TT}^{n}(Q^{2})$, compared with results from JLab E94-010 (3He) [92] (green triangles) and E97-110 (3He) [71] (blue triangles). Across all $Q^{2}$ bins, the size of the systematic uncertainty varies from $0.0275$ to $0.141$, while the relative size of the systematic to statistical uncertainty varies within $(49-123)\%$. The GDH value is shown as the short horizontal line at $\overline{I}^{n}_{TT}(0)=-0.915$. Note that earlier CLAS data [102] and the phenomenological model predictions [54, 56] are not available for this integral. See Fig. 16 caption for more details.
The resulting $\overline{\Gamma}_{1}^{n}(Q^{2})$ and $\overline{I}_{TT}^{n}(Q^{2})$ are shown in Figs. 22 and 23, respectively. They agree with previous data from E143 [107], EG1a [94, 95] and EG1b [102, 103] where the neutron information was extracted from the combined deuteron and proton data, as well as those from E94-010 [93] and E97-110 [71] for which the neutron information was extracted from 3He data. There is also agreement with the phenomenological models of Burkert and Ioffe [53, 54] and Pasechnik et al. [56] (when available) and with the $\chi$EFT predictions from Bernard et al. [59] (up to $Q^{2}\simeq 0.08$ GeV2) and, to a lesser extent, with Alarcón et al. [60]. The MAID model prediction is systematically more positive than the data. Like in the case of $\overline{\gamma}_{0}^{n}$, the calculation of [154] (not shown in the figure) agrees in sign and behavior with the $\overline{\gamma}_{0}^{n}(Q^{2})$, up to a similar normalization factor as for $\overline{\gamma}_{0}^{p}(Q^{2})$.
@@
Results on the generalized longitudinal spin polarizability on the neutron are shown in Fig. 24, along with earlier results from JLab EG1b [102, 103]), E97-110 [71] and E94-010 [93]). The EG4 neutron results generally agree with the earlier data, though there is a tension between EG4 and E97-110 for $Q^{2}\lesssim 0.1$ GeV2. Adding the EG4 deuteron and proton systematic uncertainties linearly, rather than in quadrature, does not remove the tension. It contrasts with the agreement between EG4 and E97-110 for the neutron data on $\overline{\Gamma}_{1}^{n}$ or $\overline{I}_{TT}^{n}$, suggesting the discrepancy arises from the high-$x$ (near pion threshold) contribution. The EG4 results agree with $\chi$EFT calculation from Ref. [59] up to $Q^{2}\approx 0.1$ GeV2 but not with that of Alarcón et al. [60].
Figure 24: [Color online] EG4 results on the generalized spin polarizability $\overline{\gamma}_{0}^{n}$, compared with earlier results from JLab EG1b (proton, deuteron) [102, 103] (black crossbars), E97-110 (3He) [71] (blue triangles), and E94-010 (3He) [93] (green triangles). Across all $Q^{2}$ bins, the size of the systematic uncertainty varies from $0.0158$ to $0.397$ (×$\times 10^{-4}$ fm4), while the relative size of the systematic to statistical uncertainty varies from $(40-124)\%$. Note that phenomenological model predictions [54, 56] are not available for this integral. See Fig. 16 caption for more details.
@@ VII.6 Isospin Analysis on the Moments of Spin Structure Functions
where $\omega_{D}$ is given under Eq. (38) and with the deuteron moments understood as “per nucleus”. Alternatively to using the EG4 proton and deuteron data, one can use the proton results from EG4 and the neutron results obtained from 3He data such as those from JLab E97-110 [71] to perform the isospin separation.
@@ VII.6.1 Results on the Bjorken Sum $\overline{\Gamma}_{1}^{p-n}(Q^{2})$
Figure 25: [Color online] The Bjorken sum $\overline{\Gamma}_{1}^{p-n}(Q^{2})$ from EG4 (solid red circles) and from the EG4 proton data combined with the JLab E97-110 neutron data [71] (blue triangles). Also shown are earlier data from E94-010/EG1a [42] (green triangles), CLAS EG1b [43] (black crossbars), and EG1dvcs [44] (brown open squares). The inner (outer) error bars on the data points give their statistical (total) uncertainties, while the systematic uncertainties of EG4 and EG4/E97-110 are also given by the red and the blue bands, respectively. Across all $Q^{2}$ bins, the size of the systematic uncertainty of the EG4 results (red colored band) varies from $0.0017$ to $0.0115$, while the relative size of the systematic to statistical uncertainty varies from $(58-131)\%$. The $\chi$EFT predictions from Alarcón et al. [60] and Bernard et al. [59] (the latter without error band as given in Ref. [59]) are shown along with phenomenological model predictions from Burkert-Ioffe [53, 54], Pasechnik et al. [56], MAID [113], HLFQCD [58], and Gabdrakhmanov et al. [161].
@@
Note the comparison between the $\overline{\Gamma}_{1}^{p-n}$ results obtained from EG4’s proton and deuteron data and those extracted from the EG4 proton and E97-110 neutron (3He) data. An agreement between the two different approaches was not assured since the nuclear structures of the deuteron and 3He are quite different: 3He is strongly bound compared to the deuteron, therefore more sensitive to potential nuclear modification of the proton. Our treatment of the deuteron is minimal and consists of accounting for its D-state [116, 117, 118, 119, 120], while treatment of the 3He data accounted for the nucleon effective polarizations [162]. As shown in Fig. 25, the two approaches agree well with each other. This suggests that, at least for $\overline{\Gamma}_{1}^{p-n}$, the nuclear corrections applied to the deuteron and 3He data to obtain the neutron information appear sufficient at these low $Q^{2}$ values.
@@
Both the EG4 and EG4/E97-110 results on $\overline{\Gamma}_{1}^{p-n}$ are systematically above $\chi$EFT and model predictions, except that of HLFQCD [58] and [161], which agree with the data. To allow a quantitative comparison between the data and theory, we employed a fit of the form $bQ^{2}+cQ^{4}$ to the entire world data set, based on Eq. (32[52], where $b$ is expected to be the GDH slope. Our results are shown in Table 4. The “$uncor$” uncertainty designates the point-to-point uncorrelated uncertainty. It is the quadratic sum of the statistical uncertainty and a fraction of the systematic uncertainty determined so that $\chi^{2}$/n.d.f = 1 for the best fit. The “$cor$” uncertainty is the correlated uncertainty estimated from the remaining fraction of the systematic uncertainty.
We note that our value for $b$ deviates significantly from that expected from the GDH sum rule. This is partially due to the large cancellation for the GDH sum rule in the $p-n$ combination, which is only 1/4 of the value for the individual nucleons. Experimentally, because of the nearly perfect cancellation of the large contribution from the $\Delta^{(1232)}$ resonance in the $p-n$ difference, the isovector integral is much more strongly affected by small-$x$ contributions than the other moments, with relatively large uncertainties for $g_{1}^{p}-g_{1}^{n}$. Of course, our results could also indicate a much stronger $Q^{2}$–dependence near the photon point than can be captured by a 2-parameter fit. In any case, they do not indicate a violation of the GDH sum rule since, as shown in the previous section, our results for the proton and neutron are individually consistent with the sum rule within (1 - 1.3)$\sigma$. Further discussions on the possible discrepancy between the $b$ parameter and the GDH slope can be found in Ref. [52].